# Evaluating Sparse Autoencoders for Monosemantic Representation

### v1

@@ Abstract
A key barrier to interpreting large language models is polysemanticity, where neurons activate for multiple unrelated concepts. Sparse autoencoders (SAEs) have been proposed to mitigate this issue by transforming dense activations into sparse, more interpretable features. While prior work suggests that SAEs promote monosemanticity, there has been no quantitative comparison with their base models. This paper provides the first systematic evaluation of SAEs against base models concerning monosemanticity. We introduce a fine-grained concept separability score based on the Jensen–Shannon distance, which captures how distinctly a neuron’s activation distributions vary across concepts. Using Gemma-2-2B and multiple SAE variants across five benchmarks, we show that SAEs reduce polysemanticity and achieve higher concept separability. However, greater sparsity of SAEs does not always yield better separability and often impairs downstream performance. To assess practical utility, we evaluate concept-level interventions using two strategies: full neuron masking and partial suppression. We find that, compared to base models, SAEs enable more precise concept-level control when using partial suppression. Building on this, we propose Attenuation via Posterior Probabilities (APP), a new intervention method that uses concept-conditioned activation distributions for targeted suppression. APP outperforms existing approaches in targeted concept removal.
Introduction
@@
Large language models (LLMs) have achieved remarkable performance across a wide range of natural language tasks, often matching or surpassing human-level performance (Luo et al. 2025; Achiam et al. 2023; Touvron et al. 2023; Guo et al. 2025; Eslamian and Cheng 2025). Nonetheless, understanding how these models internally represent and manipulate concepts remains a major challenge. A key obstacle is polysemanticity: the phenomenon where individual neurons respond to multiple, semantically distinct concepts rather than encoding single, interpretable features (Janiak, Mathwin, and Heimersheim 2023; Olah, Mordvintsev, and Schubert 2017; Nguyen, Yosinski, and Clune 2016). This entanglement complicates the interpretation and analysis of model behavior, posing a significant barrier to building transparent and controllable AI systems (Sharkey et al. 2025; Marshall and Kirchner 2024; Bereska and Gavves 2024).
@@
Dictionary learning via sparse autoencoders (SAEs) (Huben et al. 2024; Gao et al. 2025) has recently emerged as a promising approach to mitigating polysemanticity in neural representations. SAEs aim to transform dense activations of a desired component of the base LLM into sparse features by enforcing sparsity and encouraging each neuron to specialize in distinct, concept-specific features (Huben et al. 2024; Rajamanoharan et al. 2024a, b; Gao et al. 2025). The goal is to produce monosemantic representations, where individual neurons respond to single, well-defined concepts (Huben et al. 2024; Rajamanoharan et al. 2024a, b; Gao et al. 2025). The underlying hypothesis is intuitive: if we can force the model to use fewer neurons simultaneously, each active neuron should correspond to a more distinct and interpretable concept. Empirical studies have shown that SAEs can uncover interpretable features across domains such as vision and language, facilitating improved interpretability (Shu et al. 2025; Huben et al. 2024; Pach et al. 2025).
However, most of the evidence for the interpretability benefits of SAEs is qualitative, relying on case studies or anecdotal neuron visualizations, offering limited insight into their systematic effectiveness (Kissane et al. 2024; Li et al. 2025). While SAEs have higher computational costs and degraded downstream performance compared to their dense counterparts (Huben et al. 2024), no work has quantitatively compared their internal representations for monosemanticity: a core requirement for enhancing model interpretability. Without such comparison, the practical and theoretical value of sparsity remains poorly understood.
@@
In this work, we conduct a systematic investigation into the effectiveness of SAEs in promoting monosemanticity in the internal representations of LLMs. In contrast to prior work that evaluates SAEs in isolation, we compare their representations to those of the corresponding base models, providing the first quantitative assessment of how sparsity affects concept disentanglement. Particularly, we conduct comprehensive evaluations using Gemma-2-2B (Rivière et al. 2024) and various SAEs of different widths and sparsity levels using five benchmark datasets. We begin by quantifying polysemanticity using overlap statistics, measuring the fraction of salient neurons that respond to multiple, semantically distinct concepts. While SAEs exhibit lower polysemanticity than their base models, this overlap-based analysis remains coarse-grained: it treats all neurons that respond to multiple concepts as equally entangled, without considering how their activations vary across those concepts. In practice, a neuron may activate for several concepts, yet do so with clearly distinct activation distributions, suggesting behavior that may still be considered monosemantic. That is, monosemanticity is not solely about binary activation overlap, but rather about the separability of a neuron’s activation distributions across concepts.
@@
We formalize this view by introducing a new concept separability score, based on the Jensen–Shannon distance (Lin 1991). This fine-grained, distribution-aware metric quantifies how well a neuron’s activations separate across different concepts by measuring the distance between their activation distributions. Using this score, we find that SAEs exhibit higher concept separability than their dense counterparts. Interestingly, we also observe that increasing sparsity does not always lead to better separability, and often comes at the cost of degraded downstream performance.
@@
To further evaluate the practical utility of monosemantic representations of SAEs, we examine their effectiveness in enabling concept-level model interventions. Specifically, we assess how precisely concept-related behavior can be suppressed in SAEs compared to base models. We evaluate two intervention strategies: full neuron masking, which suppresses all activations of salient neurons associated with a target concept, and partial suppression, which intervenes in activations selectively based on their distributional association with the concept. Across SAEs and the base model, partial suppression outperforms full masking in most cases, achieving more effective suppression of the target concept while better preserving unrelated model behavior. Furthermore, our results show that SAEs support more precise and effective concept removal than their dense counterparts when applying partial suppression methods. These findings reinforce the idea that concept separability, when defined in terms of activation distributions rather than binary activation overlap, offers better model control and interpretability.
@@
Additionally, motivated by the varying separability of concept activations across neurons, we introduce Attenuation via Posterior Probabilities (APP), a new intervention method that leverages concept-conditioned activation distributions to selectively suppress target concepts with minimal side effects. Specifically, APP computes the posterior probability that a given activation corresponds to a target concept and attenuates it accordingly. Among all methods evaluated, APP achieves the strongest and most targeted concept removal across both SAEs and base models.
@@
  • We present the first quantitative analysis of monosemantic representations in SAEs relative to their base LLM.
@@
  • We propose a new intervention method, a minimally invasive concept erasure technique, outperforming existing methods in both precision and side-effect minimization.
Preliminaries
@@
Concept. A concept $c_{i}\in C$ is a semantic category assigned to each input (or its components), where $C={c_{1},\dots,c_{k}}$. For instance, sentence types (e.g., declarative, interrogative) or word-level tags (e.g., noun, verb) can serve as concepts. We focus on sentence-level concepts with one label per input.
Datasets and Models. We use five datasets: IMDB (Maas et al. 2011), AG News (Zhang, Zhao, and LeCun 2015), Emotions (Saravia et al. 2018), DBpedia (Zhang, Zhao, and LeCun 2015), and SST-2 (Socher et al. 2013). Our analysis is based on the Gemma-2-2B language model (Rivière et al. 2024) and several Sparse Autoencoders (SAEs) from GemmaScope (Lieberum et al. 2024).
@@
Analyzing Polysemanticity in SAEs
@@
A neuron is considered polysemantic when it responds to multiple, distinct concepts rather than a single, well-defined one. Several studies have demonstrated that such polysemantic behavior is common in neural networks (Elhage et al. 2022; Bau et al. 2017; Scherlis et al. 2022; Lecomte et al. 2024; Marshall and Kirchner 2024; Olah, Mordvintsev, and Schubert 2017; Nguyen, Yosinski, and Clune 2016). This polysemanticity reduces the interpretability of models, motivating the development of Sparse Autoencoders (SAEs) (Huben et al. 2024; Rajamanoharan et al. 2024a, b; Gao et al. 2025). SAEs are designed to encourage sparsity in neural activations, aiming to align each neuron with a specific, distinct concept and thereby promote monosemanticity and interpretability. In the following, we quantitatively evaluate SAEs to better understand their effectiveness in improving monosemanticity.
@@
Figure 2: Across base model and SAEs (SAE-16k, SAE-65k), neurons exhibit varying degrees of separability in their activations. Some have completely overlapping activations across concepts, others show partial or clear separation. This variability underscores the importance of using distribution-aware metrics when assessing neuron monosemanticity.
Salient Neuron Overlap
@@
As a first step, we quantify polysemanticity by measuring the overlap percentage of salient neurons (i.e., neurons with high mean activation) across concepts. Specifically, for each concept, we identify the top 50 salient neurons by mean activation. Then, the overlap percentage is computed as the intersection-over-union of these top-K sets across concepts. This metric captures the extent to which neurons are shared across concepts, reflecting shared saliency and polysemanticity. The Figure 1 compares this shared saliency for the base model and two Sparse Autoencoders (SAEs) with different latent dimensions (16k and 65k), while maintaining comparable sparsity levels (116 vs. 93 active neurons). In the DBpedia dataset, we observe that nearly 12% of the top-activated (salient) neurons in the base model are shared across all 14 concepts. On the other hand, for the IMDb and SST-2 datasets, this percentage is considerably higher as these datasets contain only two concepts, which increases the likelihood that salient neurons are shared across both. Moreover, the Figure 1 confirms that SAEs exhibit reduced conceptual overlap, suggesting less polysemanticity; however, polysemantic neurons are still present. For example, in the AG News dataset, the percentage of shared salient neurons drops from around 30% in the base model to approximately 10% in the SAE with 16k dimensions. This is still a relatively high percentage, indicating that over 10% of the most salient neurons are shared across all four concepts. Moreover, another notable observation is that the 65k-dimensional SAEs exhibit lower polysemanticity than their 16k-dimensional counterparts across all datasets. This reinforces the idea that larger SAEs have greater capacity to allocate distinct neurons to specific concepts, thereby enhancing interpretability. For additional analysis evaluating all active SAE neurons, not just the top 50, see Appendix E.
(a) SST-2 (16 k)
(b) IMDB (16 k)
(c) AG News (16 k)
(d) SST-2 (65 k)
(e) IMDB (65 k)
(f) AG News (65 k)
Figure 3: While increased sparsity generally enhances concept separability, excessive sparsity can sometimes lead to a significant drop in separability.
SAEs Activation Distributions
@@
So far, we have analyzed polysemanticity by quantifying neuron overlap based on activation frequency, specifically, how often a neuron is active or salient across different concepts. While these approaches offer useful aggregate insights, they treat neuron activation as a binary or averaged signal, overlooking the distributional characteristics of how neurons respond to each concept. In other words, a neuron might be shared across multiple concepts, but the manner in which it activates for each could vary significantly, ranging from broad, overlapping responses to distinct, well-separated patterns. To better capture this nuance, Figure 2 displays the full activation distributions of selected neurons across concept classes in the IMDB dataset, comparing the base model with two SAE variants of differing capacities. We include an analogous visualization for the AG News dataset in Appendix A.
@@
These plots reveal two key patterns: (1) consistent with prior observations (Haider et al. 2025), neuron activations in both the base model and SAEs tend to follow approximately Gaussian distributions; and (2) while some neurons are shared across multiple concepts, their activation distributions can range from highly overlapping to clearly separable. This underscores a fundamental limitation of mean-based or binary overlap measurements, which can overlook meaningful distinctions in activation behavior. To more accurately measure polysemanticity, it is therefore necessary to employ a metric that captures the full shape of activation distributions across concepts. To this end, we introduce a new concept separability score, based on the Jensen-Shannon distance, which quantifies the degree of separation between concept-specific activation distributions.
IMDB AG News Emotions DBPedia SST2
Base 0.259 0.339 0.296 0.405 0.183
@@
Concept Separability Score. To quantify how separable a neuron’s activation distributions are across $k$ concepts, let
@@
and define the mixture
$M^{(l,j)}(x)=\frac{1}{k}\sum_{i=1}^{k}f_{h^{l}_{j}\mid c_{i}}(x).$
We then define the standard Jensen–Shannon divergence (Lin 1991) as
@@
Next, we normalize to get a distance in $[0,1]$:
@@
Moreover, we assign $D_{\mathrm{JS}}=1$ whenever a neuron’s activations are all attributed to a single concept. Finally, the layer‐level separability is
@@
Table 1 reports the concept separability score $S\in[0,1]$ for layer 25, where higher values reflect more distinct activation distributions across concepts, indicating greater monosemanticity. Across all five datasets, both Sparse Autoencoders (SAEs) substantially improve separability compared to the base model. For example, on DBpedia (14 classes), the score rises from 0.405 in the base model to 0.621 and 0.680 for SAEs with 16 K and 65 K dimensions respectively, an increase of over 50%.
In most cases, increasing SAE capacity yields higher separability, consistent with the idea that more expressive latent spaces better disentangle concepts. However, two exceptions emerge. On SST2, separability actually decreases when moving from 16 K to 65 K dimensions. A similar, though smaller, decline is observed in the Emotions dataset.
Impact of SAE Sparsity
With our new separability score, we investigate two key questions: (1) Does increasing SAE sparsity (i.e., reducing the $L_{0}$ norm) consistently improve concept separability and thus, interpretability? (2) Does increasing sparsity also lead to a consistent drop in model performance? To address these questions, we present Figure 3, which illustrates how the separability score $S\in[0,1]$ (a proxy for interpretability) and the F1 score vary with increasing SAE sparsity.
Effect of Sparsity on Concept Separability. As shown in Figure 3, concept separability usually increases with greater sparsity, indicating improved interpretability. However, interestingly, at some points, excessive sparsity can degrade separability. For instance, in the SST-2 dataset (subplot (a)), the separability score drops by 18.06% when the number of active neurons ($L_{0}$ norm) is decreased from 116 to 55. A similar trend appears in the IMDB dataset (subplot (e)), with an 11.50% drop as the number of active neurons decreases from 48 to 15. A smaller decline is also observed in AG News (subplot (c)).
Effect of Sparsity on Model Performance. Increasing SAE sparsity by reducing the number of active neurons, generally makes reconstruction harder, leading to lower F1 scores in most cases. Interestingly, however, greater sparsity occasionally yields slight performance gains. For instance, in the SST-2 dataset (subplot (d), Figure 3), the F1 score improves by 16.74% as the number of active neurons decreases from 93 to 48 ($L_{0}=93$ to $L_{0}=48$).
Our analysis focused on the final transformer layer of the Gemma model (layer 25). To ensure the robustness of our findings, we repeated the experiments on layer 24 and observed similar patterns; see Appendix D.
@@
Concept Erasure
Concept erasure encompasses interventions that aim to remove a specific concept from a model’s internal representation, ideally without affecting other concepts (Dalvi et al. 2019a, b; Dai et al. 2022; Morcos et al. 2018). Formally, let $M$ be a trained model that maps an input $x$ to a concept label $M(x)=c$. An ideal erasure yields a modified model $M^{\prime}_{\mathrm{ideal}}$ satisfying:

### v2

@@ Abstract
A key barrier to interpreting large language models is polysemanticity, where neurons activate for multiple unrelated concepts. Sparse autoencoders (SAEs) have been proposed to mitigate this issue by transforming dense activations into sparse, more interpretable features. While prior work suggests that SAEs promote monosemanticity, no quantitative comparison has examined how concept activation distributions differ between SAEs and their base models. This paper provides the first systematic evaluation of SAEs against base models through activation distribution lens. We introduce a fine-grained concept separability score based on the Jensen–Shannon distance, which captures how distinctly a neuron’s activation distributions vary across concepts. Using two large language models (Gemma-2-2B and DeepSeek-R1) and multiple SAE variants across five datasets (including word-level and sentence-level), we show that SAEs reduce polysemanticity and achieve higher concept separability. To assess practical utility, we evaluate concept-level interventions using two strategies: full neuron masking and partial suppression. We find that, compared to base models, SAEs enable more precise concept-level control when using partial suppression. Building on this, we propose Attenuation via Posterior Probabilities (APP), a new intervention method that uses concept-conditioned activation distributions for targeted suppression. APP achieves the smallest perplexity increase while remaining highly effective at concept removal  [^1] Code will be released publicly after paper acceptance..
Evaluating Sparse Autoencoders for Monosemantic Representation
Moghis Fereidouni, Muhammad Umair Haider, Peizhong Ju, A.B. Siddique, University of Kentucky, moghis.fereidouni@uky.edu, muhammadumairhaider@uky.edu, peizhong.ju@uky.edu, siddique@cs.uky.edu
1 Introduction
@@
Large language models (LLMs) have achieved remarkable performance across a wide range of natural language tasks, often matching or surpassing human-level performance Luo et al. (2025); Achiam et al. (2023); Touvron et al. (2023); Guo et al. (2025); Eslamian and Cheng (2025). Nonetheless, understanding how these models internally represent and manipulate concepts remains a major challenge. A key obstacle is polysemanticity; the phenomenon where individual neurons respond to multiple, semantically distinct concepts rather than encoding single, interpretable features Janiak et al. (2023); Olah et al. (2017); Nguyen et al. (2016). This entanglement complicates the interpretation and analysis of model behavior, posing a significant barrier to building transparent and controllable AI systems Sharkey et al. (2025); Marshall and Kirchner (2024); Bereska and Gavves (2024).
@@
Dictionary learning via sparse autoencoders (SAEs) Huben et al. (2024); Gao et al. (2025) has recently emerged as a promising approach to mitigating polysemanticity in neural representations. SAEs aim to transform dense activations of a desired component of the base LLM into sparse features by enforcing sparsity and encouraging each neuron to specialize in distinct, concept-specific features Huben et al. (2024); Rajamanoharan et al. (2024a, b); Gao et al. (2025). The goal is to produce monosemantic representations, where individual neurons respond to single, well-defined concepts Huben et al. (2024); Rajamanoharan et al. (2024a, b); Gao et al. (2025). The underlying hypothesis is intuitive; if we can force the model to use fewer neurons simultaneously, each active neuron should correspond to a more distinct and interpretable concept. Empirical studies have shown that SAEs can uncover interpretable features across domains such as vision and language, facilitating improved interpretability Shu et al. (2025); Huben et al. (2024); Pach et al. (2025).
Most existing evaluations of SAE interpretability are qualitative, relying on case studies or anecdotal neuron visualizations that provide limited systematic insight Kissane et al. (2024); Li et al. (2025). The remaining quantitative efforts mainly examine whether neurons are active for given concepts, rather than capturing the full distributional structure of concept activations across neurons Minegishi et al. (2025); Karvonen et al. (2025).
@@
In this work, we conduct a systematic investigation into the effectiveness of SAEs in promoting monosemanticity in the internal representations of LLMs through the distributional lens. Particularly, we conduct comprehensive evaluations using two large language models (Gemma-2-2B Rivière et al. (2024) and DeepSeek-R1 Guo et al. (2025)) and various SAEs of different widths and sparsity levels on five benchmark datasets, including both word-level (POS tagging and NER) and sentence-level (e.g., AG News) tasks. We begin by quantifying polysemanticity using overlap statistics, measuring the fraction of salient neurons that respond to multiple, semantically distinct concepts. While SAEs exhibit lower polysemanticity than their base models, this overlap-based analysis remains coarse-grained; it treats all neurons that respond to multiple concepts as equally entangled, without considering how their activations vary across those concepts. In practice, a neuron may activate for several concepts, yet do so with clearly distinct activation distributions, suggesting behavior that may still be considered monosemantic. That is, monosemanticity is not solely about binary activation overlap, but rather about the separability of a neuron’s activation distributions across concepts.
@@
We formalize this view by introducing a new concept separability score, based on the Jensen–Shannon distance Lin (1991). This fine-grained, distribution-aware metric quantifies how well a neuron’s activations separate across different concepts by measuring the distance between their activation distributions. Using this score, we find that SAEs exhibit higher concept separability than their dense counterparts.
@@
To further evaluate the practical utility of monosemantic representations of SAEs, we examine their effectiveness in enabling concept-level model interventions. Specifically, we assess how precisely concept-related behavior can be suppressed in SAEs compared to base models. We evaluate two intervention strategies; full neuron masking, which suppresses all activations of salient neurons associated with a target concept, and partial suppression, which intervenes in activations selectively based on their distributional association with the concept. Across SAEs and the base model, partial suppression outperforms full masking in most cases, achieving more effective suppression of the target concept while better preserving unrelated model behavior. Furthermore, our results show that SAEs support more precise and effective concept removal than their dense counterparts, especially when applying partial suppression methods. These findings reinforce the idea that concept separability, when defined in terms of activation distributions rather than binary activation overlap, offers better model control and interpretability.
@@
Additionally, motivated by the varying separability of concept activations across neurons, we introduce Attenuation via Posterior Probabilities (APP), a new intervention method that leverages concept-conditioned activation distributions to selectively suppress target concepts with minimal side effects. Specifically, APP computes the posterior probability that a given activation corresponds to a target concept and attenuates it accordingly. Among all methods evaluated, APP achieves the smallest degradation in language modeling quality (lowest perplexity increase) while remaining highly competitive with other baselines in targeted concept removal across both SAEs and base models.
@@
  • We present the first quantitative analysis of monosemantic representations in SAEs relative to their base LLM through distributional lens.
@@
  • We propose a new intervention method, which is the least invasive concept erasure technique and highly competitive with existing methods in removing the targeted concept.
2 Preliminaries
@@
Concept. A concept $c_{i}\in C$ is a semantic category assigned to each input (or its components), where $C={c_{1},\dots,c_{k}}$. For instance, sentence-level types (e.g., declarative, interrogative) or word-level tags (e.g., noun, verb) can serve as concepts. In this work, we focus both on sentence-level and word-level concepts.
Datasets and Models. We use five datasets: Part-of-Speech Tagging (POS) Pasini et al. (2021), Named Entity Recognition (NER) Ding et al. (2021), AG News Zhang et al. (2015), Emotions Saravia et al. (2018), and DBpedia Zhang et al. (2015). Our analysis is conducted using two large language models: (1) the Gemma-2-2B model Rivière et al. (2024) along with its corresponding Sparse Autoencoders (SAEs) from GemmaScope Lieberum et al. (2024), and (2) the DeepSeek-R1 model Guo et al. (2025) with its associated SAE from LlamaScope He et al. (2024).
@@
3 Analyzing Polysemanticity in SAEs
@@
A neuron is considered polysemantic when it responds to multiple, distinct concepts rather than a single, well-defined one. Several studies have demonstrated that such polysemantic behavior is common in neural networks Elhage et al. (2022); Bau et al. (2017); Scherlis et al. (2022); Lecomte et al. (2024); Marshall and Kirchner (2024); Olah et al. (2017); Nguyen et al. (2016). This polysemanticity reduces the interpretability of models, motivating the development of Sparse Autoencoders (SAEs) Huben et al. (2024); Rajamanoharan et al. (2024a, b); Gao et al. (2025). SAEs are designed to encourage sparsity in neural activations, aiming to align each neuron with a specific, distinct concept and thereby promote monosemanticity and interpretability. In the following, we evaluate SAEs to better understand their effectiveness in improving monosemanticity.
@@
Figure 2: Across base model and SAEs (SAE-16k, SAE-65k), neurons exhibit varying degrees of separability in their activations. Some have completely overlapping activations across concepts, others show partial or clear separation. This variability underscores the importance of using distribution-aware metrics when assessing monosemanticity.
3.1 Salient Neuron Overlap
@@
As a first step, we quantify polysemanticity by measuring the overlap percentage of salient neurons (i.e., neurons with high mean activation) across concepts. Specifically, for each concept, we identify the top 80 salient neurons by mean activation. Then, the overlap percentage is computed as the intersection-over-union of these top-K sets across concepts. This metric captures the extent to which neurons are shared across concepts, reflecting shared saliency and polysemanticity. The Figure 1 compares this shared saliency for the base model and two Sparse Autoencoders (SAEs) with different latent dimensions (16k and 65k), while maintaining comparable sparsity levels (116 vs. 93 active neurons). In the DBpedia dataset, for instance, we observe that nearly 9% of the top-activated (salient) neurons in the base model are shared across all 14 concepts. Moreover, the Figure 1 confirms that SAEs exhibit reduced conceptual overlap, suggesting less polysemanticity; however, polysemantic neurons are still present. For example, in the POS dataset, the percentage of shared salient neurons drops from around 38% in the base model to approximately 18% in the SAE with 16k dimensions. This is still a relatively high percentage, indicating that over 18% of the most salient neurons are shared across all four concepts. Moreover, another notable observation is that the 65k-dimensional SAEs exhibit lower polysemanticity than their 16k-dimensional counterparts across all datasets. This reinforces the idea that larger SAEs have greater capacity to allocate distinct neurons to specific concepts, thereby enhancing interpretability. For additional analysis evaluating all active SAE neurons, not just the top 80, see Appendix E.
3.2 SAEs Activation Distributions
@@
So far, we have analyzed polysemanticity by quantifying neuron overlap based on activation frequency, specifically, how often a neuron is active or salient across different concepts. While these approaches offer useful aggregate insights, they treat neuron activation as a binary or averaged signal, overlooking the distributional characteristics of how neurons respond to each concept. In other words, a neuron might be shared across multiple concepts, but the manner in which it activates for each could vary significantly, ranging from broad, overlapping responses to distinct, well-separated patterns. To better capture this nuance, Figure 2 displays the full activation distributions of selected neurons across concept classes in the AG News dataset, comparing the base model with two SAE variants of differing capacities. We include an analogous visualization for the NER dataset in Appendix B.
@@
These plots reveal two key patterns: (1) consistent with prior observations Haider et al. (2025), neuron activations in both the base model and SAEs tend to follow approximately Gaussian distributions; and (2) while some neurons are shared across multiple concepts, their activation distributions can range from highly overlapping to clearly separable. This underscores a fundamental limitation of mean-based or binary overlap measurements, which can overlook meaningful distinctions in activation behavior. To more accurately measure polysemanticity, it is therefore necessary to employ a metric that captures the full shape of activation distributions across concepts. To this end, we introduce a new concept separability score, based on the Jensen-Shannon distance, which quantifies the degree of separation between concept-specific activation distributions.
POS AG News Emotions DBpedia NER
Base 0.343 0.366 0.322 0.405 0.308
@@
Concept Separability Score. To quantify how separable a neuron’s activation distributions are across $k$ concepts, we first define the probability density function as:
@@
with this density function at hand, we define mixture as $M^{(l,j)}(x)=\frac{1}{k}\sum_{i=1}^{k}f_{h^{l}_{j}\mid c_{i}}(x)$,
and then compute the generalized Jensen–Shannon divergence Lin (1991) as
@@
Next, we take the square root of the JSD and normalize it to obtain a proper distance metric bounded within $[0,1]$:
@@
Moreover, we assign $D_{\mathrm{JS}}=1$ whenever a neuron’s activations are all attributed to a single concept. Finally, the layer‐level separability score is
@@
Table 1 reports the concept separability score $S\in[0,1]$ for Gemma, where higher values indicate more distinct activation distributions across concepts, indicating higher monosemanticity (Full JS scores for all settings and both LLMs are in Appendix G). Across all five datasets, SAEs substantially improve separability over the base model. For instance, on DBpedia (14 classes), the score increases from 0.405 in the base model to 0.621 and 0.680 for the 16K- and 65K-dimensional SAEs, respectively, an increase of over 50%. Furthermore, higher-capacity SAEs consistently yield greater separability, supporting the view that more expressive latent spaces better disentangle concepts.
@@
4 Concept Erasure
Concept erasure encompasses interventions that aim to remove a specific concept from a model’s internal representation, ideally without affecting other concepts Dalvi et al. (2019a, b); Dai et al. (2022); Morcos et al. (2018). Formally, let $M$ be a trained model that maps an input $x$ to a concept label $M(x)=c$. An ideal erasure yields a modified model $M^{\prime}_{\mathrm{ideal}}$ satisfying: